Company’s relevant range of production is 7,500 units to 12,500 units. When it produces and sells 10,000 units, its average costs per unit are as follows:

Average Cost Per Unit
Direct materials $ 5.90
Direct labor $ 3.40
Variable manufacturing overhead $ 1.60
Fixed manufacturing overhead $ 4.00
Fixed selling expense $ 2.90
Fixed administrative expense $ 2.20
Sales commissions $ 1.20
Variable administrative expense $ 0.45

14. If 10,000 units are produced, what are the total amounts of direct and indirect manufacturing costs incurred to support this level of production? (Do not round intermediate calculations.)

15. What incremental manufacturing cost will Martinez incur if it increases production from 10,000 to 10,001 units? (Round your answer to 2 decimal places.)
Answer #1

14.

Direct Material $     59,000.00 =10000*5.90
Direct Labor $     34,000.00 =10000*3.4
Total Direct Manufacturing cost $    93,000.00
Variable Manufacturing Overhead $     16,000.00 =10000*1.6
Fixed Manufacturing Overhead $     40,000.00 =10000*4
Total Indirect Manufacturing cost $    56,000.00

15.
Incremental Manufacturing cost = $5.90 + 3.40 + 1.60 = $10.90 per unit